# On the representation of stack operators by mathematical morphology

### v1

@@ 5 Representation of stack W-operators
The representation of set and lattice operators by mathematical morphology is a classical research topic, and results about the representation of increasing set operators [7], translation-invariant set operators [1], and general lattice operators [2] have been established. In this section we deduce the representation of stack W-operators by a characteristic function, a kernel and a basis, and establish isomorphisms with the respective representation of the associated set W-operator, that are illustrated in Figure 2.
[…]
@@ 5.1 Representation by characteristic function
Define $\mathcal{P}(W)\coloneqq\{0,1\}^{\chi(W)}$ and $\mathcal{P}_{M}(W)\coloneqq M^{\chi(W)}$. For each $\tilde{\psi}\in\Psi^{W}$ and $\psi\in\Sigma_{S}^{W}$, define the characteristic functions $\phi_{\psi}:\mathcal{P}_{M}(W)\mapsto M$ and $\tilde{\phi}_{\tilde{\psi}}:\mathcal{P}(W)\mapsto\{0,1\}$ as $\phi_{\psi}(f)=\psi(f)(o)$ and $\tilde{\phi}_{\tilde{\psi}}(X)=\tilde{\psi}(X)(o)$ for $f\in\mathcal{P}_{M}(W)$ and $X\in\mathcal{P}(W)$, recalling that $o$ is the zero element of $(E,+)$. Define $\Phi_{S}^{W}\coloneqq\{\phi_{\psi}:\psi\in\Sigma_{S}^{W}\}\subsetneq M^{\mathcal{P}_{M}(W)}$ and $\Phi^{W}\coloneqq\{0,1\}^{\mathcal{P}(W)}$ as the space of characteristic functions of stack and set W-operators, respectively, and consider in them the point-wise partial orders:
@@
It follows from Propositions 2 and 5 that there exists a lattice isomorphism $\mathcal{G}:\Phi^{W}\mapsto\Phi_{S}^{W}$ given, for example, by $\mathcal{G}=\mathcal{C}_{S}\circ\mathcal{F}\circ\mathcal{C}^{-1}$. Next lemma gives explicit formulas for $\mathcal{G}$ and $\mathcal{G}^{-1}$ that can be obtained by explicitly computing $\mathcal{C}_{S}\circ\mathcal{F}\circ\mathcal{C}^{-1}$ and its inverse.
@@ 5.2 Kernel representation
for $\tilde{\phi}\in\Phi^{W},\phi\in\Phi_{S}^{W}$ and $t\in M$, in which $\kappa_{S}^{W}\coloneqq\{\mathcal{K}_{S}(\phi):\phi\in\Phi_{S}^{W}\}\subsetneq\kappa_{M}^{W}$. The next proposition, that is direct from the definition of $\mathcal{K}$ and $\mathcal{K}_{S}$, shows that these are indeed lattice isomorphisms. The inverse of these mappings are $\mathcal{K}^{-1}(\mathscr{X})(X)=\mathds{1}\{X\in\mathscr{X}\}$ and $\mathcal{K}_{S}^{-1}(\mathscr{F})(f)=\max\{t\in M:f\in\mathscr{F}(t)\}$.
@@ 5.3 Basis representation
For $f_{1},f_{2}\in\mathcal{P}_{M}(W)$ and $X_{1},X_{2}\in\mathcal{P}(W)$, the intervals with limits $f_{1},f_{2}$ and $X_{1},X_{2}$ are the sets $[f_{1},f_{2}]=\{f\in\mathcal{P}_{M}(W):f_{1}\leq f\leq f_{2}\}$ and $[X_{1},X_{2}]=\{X\in\mathcal{P}(W):X_{1}\leq X\leq X_{2}\}$, respectively, which are empty if $f_{1}\nleq f_{2}$ or $X_{1}\nleq X_{2}$. For collections $\mathscr{X}\in\mathcal{P}(\mathcal{P}(W))$ and $\mathscr{G}\in\mathcal{P}(\mathcal{P}_{M}(W))$, define the respective maximal collection as formed by the elements in it that are not lesser than any other one: $\textbf{\text{Max}}(\mathscr{X})=\{X\in\mathscr{X}:Y\in\mathscr{X},X\leq Y\implies X=Y\}$ and $\textbf{\text{Max}}(\mathscr{G})=\{f\in\mathscr{G}:g\in\mathscr{G},f\leq g\implies f=g\}$.
Sets and stack W-operators can be uniquely represented by the maximal intervals contained in their kernel. Formally, define by $\Pi^{W}\subseteq\mathcal{P}(\mathcal{P}(W))$ the family of collections of maximal intervals in $\mathcal{P}(W)$ and by $\Pi_{S}^{W}\subseteq\mathcal{P}(\mathcal{P}_{M}(W))$ the family of collections of maximal intervals in $\mathcal{P}_{M}(W)$. Consider in these collections the partial orders
@@
In particular, the isomorphism $\mathcal{I}$ takes basis of set operators that have one interval to basis of stack operators that have one interval for every $t$. This simple result is key to prove the main result of this paper (ref. Theorem 6.1). This result is proved in Appendix 0.A.
@@
6 Restricted classes of stack operators and learning of stack operators
@@
The isomorphisms depicted in Figure 2 together with Theorem 6.1 imply that a stack W-operator can be designed by designing a set operator that performs the desired transformation on sets, what actually reduces to designing a binary function in $\mathcal{P}(W)$. Prior information about properties that the optimal operator satisfy can be used to define a restricted class of set operators that, due to Theorem 6.1, will generate a restricted class of stack operators with the same properties, from which an operator can be learned from data. In particular, it is possible to adapt the loss function considered to train discrete morphological neural networks (DMNN) [9], in particular unrestricted DMNN [10], to learn a characteristic set W-operator by actually minimising the error associated with the respective stack operator. Also, a deep neural network might also be considered to represent the characteristic function of the set W-operator and be trained by minimising the error associated with the respective stack operator. We will investigate these learning methods in future studies.
7 Final remarks
This paper defined the stack operators, that are those that commute in average with cross-sectioning, and showed that they can be represented by a characteristic set operator, being equivalent to applying the set operator to the cross-sections of the image and summing. In particular, stack operators are 1-Lipschitz extensions of set operators and also the generalisation of stack filters, for which the characteristic set operator is increasing. Isomorphisms between the characteristic function, kernel and basis representation of stack W-operators and set W-operators were deduced and we showed that many properties of the set W-operators are preserved in the respective stack W-operators. The perspectives of learning stack operators by learning the characteristic set operator via DMNN and deep neural networks was briefly discussed.
There are many topics for future studies based on our representation results about stack operators. First, it is necessary to generalise the results for stack-operators that are not W-operators, what should be straightforward. Furthermore, it would be interesting to study the representation of stack operators with other kinds of invariance, such as group equivariant operators [11] and scale invariant operators [12]. Moreover, as discussed in Section 6, it is necessary to develop learning methods for stack operators.
@@ 7 Final remarks
Given the simplicity of the representation of stack W-operators, that are equivalent to Boolean functions in $\mathcal{P}(W)$, they could be the preferred choice when a solving a practical image transformation problem. However, it is necessary to characterise what problems they can solve, that is, what image operators they can approximate. Therefore, the main topic of future research is to characterise the stack operators within the family of 1-Lipschitz extensions of set operators by investigating if they have universal approximating properties, i.e., by characterising spaces in which the stack operators are dense. Stack operators will be suitable to solve a problem when the optimal operator is in such a space.
@@ Acknowledgements
This work was partially funded by grant #22/06211-2 São Paulo Research Foundation (FAPESP).
@@ Appendix 0.A Proof of results
Fix $f,g\in M^{E}$ and $\psi\in\Sigma_{S}$, and denote $\tilde{\psi}\coloneqq\mathcal{F}^{-1}(\psi)$. Define $D(f,g)=\{t\in\{1,\dots,m\}:T_{t}[f]\neq T_{t}[g]\}$ and observe that $\lVert\psi(f)-\psi(g)\rVert_{\infty}\leq\sum_{t\in D(f,g)}\lVert(\tilde{\psi}(T_{t}[f])-\tilde{\psi}(T_{t}[g]))\rVert_{\infty}\leq|D(f,g)|$. We will show that $|D(f,g)|\leq\lVert f-g\rVert_{1}$. For each $t\in D(f,g)$ there exists $x_{t}\in E$ such that $T_{t}[f](x_{t})\neq T_{t}[g](x_{t})$. Fix a set $\{x_{t}:t\in D(f,g)\}$ that satisfies $x_{t}=x_{t+1}=\cdots=x_{t+s-1}$ whenever $T_{t-1}[f](x_{t})=T_{t-1}[g](x_{t})$, $T_{t}[f](x_{t})\neq T_{t}[g](x_{t})$ and $|f(x_{t})-g(x_{t})|=s$. It follows that $|D(f,g)|=\sum_{\{x_{t}:t\in D(f,g)\}}|f(x_{t})-g(x_{t})|\leq\lVert f-g\rVert_{1}$. $\blacksquare$
@@
=$=\left\{\sum_{s=1}^{m}X_{s}:X_{s}\in\mathcal{P}(W),X_{m}\leq\cdots\leq X_{1},\sum_{s=1}^{m}\mathds{1}\{X_{s}\in\mathscr{X}\}\geq t\right\}.$
@@
For each $\psi\in\Sigma_{S}^{W}$ denote $\tilde{\psi}\coloneqq\mathcal{F}^{-1}(\psi)$ and $\tilde{\phi}\coloneqq\mathcal{C}(\tilde{\psi}$), and for each $\tilde{\psi}\in\Psi^{W}$ denote $\psi\coloneqq\mathcal{F}(\tilde{\psi})$ and $\phi\coloneqq\mathcal{C}_{S}(\psi)$.
(1) Increasing: Direct from Proposition 4.
(2) Decreasing: Analogous to increasing.
(3) Extensive: Fix $\tilde{\psi}\in\Psi^{W}$ extensive and $f\in\mathcal{P}_{M}(W)$. Then, $\psi(f)=\sum_{t=1}^{m}\tilde{\psi}(T_{t}[f])\geq\sum_{t=1}^{m}T_{t}[f]=f$ in which the inequality follows since $\tilde{\psi}$ is extensive and the last equality is due to Proposition 1 (i). Fix $\psi\in\Sigma_{S}^{W}$ extensive and $X\in\mathcal{P}(W)$. Then, $\tilde{\psi}(X)=\frac{1}{m}\psi(mX)\geq\frac{1}{m}mX=X$.
(4) Anti-extensive: Analogous to extensive.
(5) Erosion: Set and image operators are erosions if, and only if, their kernel is formed by only one interval of form $[X,W]$ and $[f_{t},mW]$ for all $t\geq 1$, respectively. On the one hand, it follows from Corollary 2 that if $\tilde{\psi}$ is an erosion with kernel $[X,W]$ then the kernel of $\psi$ equals $[tX,W]$ for all $t\geq 1$ so it is an erosion. On the other hand, if $\psi$ is an erosion then $\mathcal{K}_{S}(\psi)(m)=[f_{m},mW]$ and $\mathcal{K}(\tilde{\psi})=\{X\in\mathcal{P}(W):f_{m}\leq mX\}$. It is clear that $f_{m}\leq mX$ if, and only if, $T_{1}[f_{m}]\leq X$ and therefore $\mathcal{K}(\tilde{\psi})=[T_{1}[f_{m}],W]$, and $\tilde{\psi}$ is an erosion.
(6) Dilation: Analogous to erosion by considering the dual partial orders since dilations according to a partial order are erosions according to the dual partial order.
(7) Anti-dilation: Anti-dilations are operators such that $\psi(f\vee g)=\psi(f)\wedge\psi(g)$ and $\tilde{\psi}(X\vee Y)=\tilde{\psi}(X)\wedge\tilde{\psi}(Y)$. We recall that the kernel of an anti-dilation is formed by a unique interval of form $[\emptyset,Y]$, for set W-operators, and $[\emptyset,g_{t}]$ for all $t\geq 1$, for image operators. A deduction analogous to that of erosions follow by Corollary 2.
@@
(10) Inf-generating: An inf-generating operator is the join between a dilation and an anti-erosion (ref. [2]). Analogous to sup-generating by considering the dual partial orders since sup-generating operators according to a partial order are inf-generating according to the dual partial order.

### v2

@@ 4 Representation of stack W-operators
The representation of set and lattice operators by mathematical morphology is a classical research topic, and results about the representation of increasing set operators [9], translation-invariant set operators [1], and general lattice operators [2] have been established. In this section, we deduce the representation of stack W-operators by a characteristic function, a kernel, and a basis, and establish isomorphisms with the respective representation of the associated set W-operator, that are illustrated in Figure 2.
$\Psi^{W}$$\Phi^{W}$$\kappa^{W}$$\Pi^{W}$$\Sigma_{S}^{W}$$\Phi_{S}^{W}$$\kappa^{W}_{S}$$(\Pi^{W}_{S})^{M}$$\mathcal{C}$$\mathcal{C}^{-1}$$\mathcal{K}$$\mathcal{K}^{-1}\mathcal{B}$$\mathcal{B}^{-1}$$\mathcal{C}_{S}$$\mathcal{C}_{S}^{-1}$$\mathcal{K}_{S}$$\mathcal{K}_{S}^{-1}$$\mathcal{B}_{S}$$\mathcal{B}^{-1}_{S}\mathcal{F}$$\mathcal{F}^{-1}$$\mathcal{G}$$\mathcal{G}^{-1}\mathcal{H}$$\mathcal{H}^{-1}\mathcal{I}$$\mathcal{I}^{-1}$
@@
There are at least two benefits of deducing representations for stack operators. On the one hand, a representation by the combination of simpler operators may yield a more efficient algorithm to compute the operator. This benefit is evident in the representation of the example operator $\psi_{1}$ by its characteristic set operator. This representation, that is a consequence of isomorphism $\mathcal{F}$, allows computing $\psi_{1}$ by applying erosions and dilations in sequence to the image cross-sections and then adding, what might be more efficient than other representations. Furthermore, if the basis (see Section 4.3) of the alternating-sequential set operator $\tilde{\psi}_{1}$ was computed, then the operator could be implemented by computing $\tilde{\psi}_{1}$ as the supremum of erosions [^5] The fact that increasing set operators can be represented as the supremum of erosions has been proved in [8]., what is more computationally efficient. This representation is a consequence of the isomorphism $\mathcal{F}\circ\mathcal{C}^{-1}\circ\mathcal{K}^{-1}\circ\mathcal{B}^{-1}$ that maps the basis of the characteristic set operator to the stack operator, as will be proved in the next sections. We also note that certain representations may yield more efficient algorithms for learning operators from data; see for example the algorithm developed in [13, 14] to learn the composition of set W-operators from data, relying on the representation via the characteristic function (see Section 4.1).
On the other hand, considering the main result of this paper (i.e. Theorem 5.1), since stack operators inherit properties of the characteristic set operators, in order to design stack operators with desired properties, it is enough to do so on the space of set operators, and it may be easier to enforce desired properties on set operators by considering suitable representations for them. For example, via the basis representation, it is easy to enforce the increasing property by representing the operator as the supremum of erosions. With this representation, it remains to design the structuring elements of these erosions, what can be done manually or by learning from data. Likewise, other restrictions may be considered on the basis of the characteristic set operator to yield desired properties, and the intervals of the basis (see Section 4.3) may be designed under these constraints (see Section 5 for more details).
@@ 4.1 Representation by characteristic function
Define $\mathcal{P}(W)\coloneqq\{0,1\}^{\chi(W)}$ and $\mathcal{P}_{M}(W)\coloneqq M^{\chi(W)}$ as the collection of binary and grey-scale patches with support in $\chi(W)$, respectively. For each $\tilde{\psi}\in\Psi^{W}$ and $\psi\in\Sigma_{S}^{W}$, define the characteristic functions $\phi_{\psi}:\mathcal{P}_{M}(W)\mapsto M$ and $\tilde{\phi}_{\tilde{\psi}}:\mathcal{P}(W)\mapsto\{0,1\}$ as $\phi_{\psi}(f)=\psi(f)(o)$ and $\tilde{\phi}_{\tilde{\psi}}(X)=\tilde{\psi}(X)(o)$ for $f\in\mathcal{P}_{M}(W)$ and $X\in\mathcal{P}(W)$, recalling that $o$ is the zero element of $(E,+)$. The characteristic function determines the value of each pixel $x\in E$ in the transformed image from the values of the pixels of the original image in its neighbourhood $x+W$.
Define $\Phi_{S}^{W}\coloneqq\{\phi_{\psi}:\psi\in\Sigma_{S}^{W}\}\subsetneq M^{\mathcal{P}_{M}(W)}$ and $\Phi^{W}\coloneqq\{0,1\}^{\mathcal{P}(W)}$ as the space of characteristic functions of stack and set W-operators, respectively, and consider in them the point-wise partial orders:
@@
It follows from Propositions 1 and 4 that there exists a lattice isomorphism $\mathcal{G}:\Phi^{W}\mapsto\Phi_{S}^{W}$ given, for example, by $\mathcal{G}=\mathcal{C}_{S}\circ\mathcal{F}\circ\mathcal{C}^{-1}$. Next lemma gives explicit formulas for $\mathcal{G}$ and $\mathcal{G}^{-1}$ that can be obtained by explicitly computing $\mathcal{C}_{S}\circ\mathcal{F}\circ\mathcal{C}^{-1}$ and its inverse.
@@ 4.2 Kernel representation
for $\tilde{\phi}\in\Phi^{W},\phi\in\Phi_{S}^{W}$ and $t\in M$, in which $\kappa_{S}^{W}\coloneqq\{\mathcal{K}_{S}(\phi):\phi\in\Phi_{S}^{W}\}\subsetneq\kappa_{M}^{W}$. This definition of kernel is due to [1, 2].
The next proposition, that is direct from the definition of $\mathcal{K}$ and $\mathcal{K}_{S}$, shows that these are indeed lattice isomorphisms. The inverse of these mappings are $\mathcal{K}^{-1}(\mathscr{X})(X)=\mathds{1}\{X\in\mathscr{X}\}$ and $\mathcal{K}_{S}^{-1}(\mathscr{F})(f)=\max\{t\in M:f\in\mathscr{F}(t)\}$.
@@ 4.3 Basis representation
For $f_{1},f_{2}\in\mathcal{P}_{M}(W)$ and $X_{1},X_{2}\in\mathcal{P}(W)$, the intervals with limits $f_{1},f_{2}$ and $X_{1},X_{2}$ are the sets $f_{1},f_{2}]=\{f\in\mathcal{P}_{M}(W):f_{1}\leq f\leq f_{2}\}$ and $X_{1},X_{2}]=\{X\in\mathcal{P}(W):X_{1}\leq X\leq X_{2}\}$, respectively, which are empty if $f_{1}\nleq f_{2}$ or $X_{1}\nleq X_{2}$. For collections $\mathscr{X}\in\mathcal{P}(\mathcal{P}(W))$ and $\mathscr{G}\in\mathcal{P}(\mathcal{P}_{M}(W))$, define the respective maximal collection as formed by the elements in it that are not lesser than any other one: $\textbf{\text{Max}}(\mathscr{X})=\{X\in\mathscr{X}:Y\in\mathscr{X},X\leq Y\implies X=Y\}$ and $\textbf{\text{Max}}(\mathscr{G})=\{f\in\mathscr{G}:g\in\mathscr{G},f\leq g\implies f=g\}$.
Since set and stack W-operators can be uniquely represented by their kernel, they can be represented by any collection of sets that covers their kernel [^6] In the case of grey-scale image operators, there should actually be a collection of sets for each $t$, that covers the kernel applied at a point $t$.. A way of covering a subset of a lattice, is by a collection of intervals, and such a cover with the minimum number of intervals is that of maximal intervals. The basis of an operator was proposed by [1, 2] based on the idea of covering the kernel (applied at point $t$ in the grey-scale case) with maximal intervals.
Formally, define by $\Pi^{W}\subseteq\mathcal{P}(\mathcal{P}(W))$ the family of collections of maximal intervals in $\mathcal{P}(W)$ and by $\Pi_{S}^{W}\subseteq\mathcal{P}(\mathcal{P}_{M}(W))$ the family of collections of maximal intervals in $\mathcal{P}_{M}(W)$. Consider in these collections the partial orders
@@
In particular, the isomorphism $\mathcal{I}$ takes basis of set operators that have one interval to basis of stack operators that have one interval for every $t$. This simple result is key to proving Theorem 5.1. This result is proved in Appendix 0.A.
@@
5 Restricted classes of stack operators and learning from data
@@
The isomorphisms depicted in Figure 2 together with Theorem 5.1 imply that a stack W-operator can be designed by designing a set operator that performs the desired transformation on sets, what actually reduces to designing a binary function in $\mathcal{P}(W)$. This is what has been done to design the stack operators in Figure 1.
Prior information about properties that the optimal stack operator satisfies can be used to define a restricted class of set operators that, due to Theorem 5.1, will generate a restricted class of stack operators with the same properties, from which an operator can be learned from data. For example, applying the isomorphism $\mathcal{F}\circ\mathcal{C}^{-1}\circ\mathcal{K}^{-1}\circ\mathcal{B}^{-1}$ to the basis $I=\{[A_{i},B_{i}]:i=1,\dots,n\}\in\Pi^{W}$ of a set operator we get the stack operator given by
$\psi_{I}(f)=\sum_{t=1}^{m}\bigvee_{i=1}^{n}\lambda_{[A_{i},B_{i}]}(T_{t}[f]),$ $f\in M^{E}$ (4)
in which $\lambda_{[A_{i},B_{i}]}$ is the set operator with characteristic function $\tilde{\phi}_{[A_{i},B_{i}]}(X)=1$ if $A_{i}\leq X\leq B_{i}$ and equal to zero otherwise for $X\in\mathcal{P}(W)$, that is the operator with basis $\{[A_{i},B_{i}]\}$. This decomposition is based on the sup-generating representation of set operators developed by [1] and $\lambda_{[A_{i},B_{i}]}$ is called a sup-generating operator.
If pairs of input and target images $(f_{j},g_{j}),j=1,\dots,N$, are available, then, based on (4), a stack operator may be learned from data by minimising the loss $L_{N}(I)=\frac{1}{N}\sum_{j=1}^{N}\lVert\psi_{I}(f_{j})-g_{j}\rVert_{1}$ over $I$ in a subset of $\Pi^{W}$ that depends on prior knowledge about the target operator, in which the size $n$ of the basis in (4) controls the complexity of the operator. For example, to learn a stack filter, we can consider only basis $I$ with $B_{i}=W$ for all [^7] It follows from the representation deduced in [8] that increasing operators have basis with intervals of form $A_{i},W]$. $i=1,\dots,n$. Also, to learn extensive stack operators we can consider only basis which contain the interval $[A,W]$ with $A(o)=1$ and $A(x)=0$ for $x\neq o$, what guarantees that $f(x)\leq\psi_{I}(f)(x)$ for all $x\in E$. Other properties may be enforced on $\psi_{I}$ by considering an inf-generating representation, under which dilations, decreasing and anti-extensive operators have simplified representations (see [1] for more details).
The minimisation of $L_{N}(I)$ can be carried out by the lattice descent algorithm proposed for training discrete morphological neural networks (DMNN) in [11] yielding a method that is an extension of DMNN for grey-scale image operators. By dropping the assumption that the intervals in $I$ are maximal, the representation (4) is actually an overparametrisation of the operator, since there are more than one collection of intervals representing the same operator, that are all collections that cover its kernel, but the optimisation problem becomes more efficient (see [12] for more details). If a characteristic function representation for the set operator was considered instead, then, relying on the isomorphism $\mathcal{F}\circ\mathcal{C}^{-1}$, the algorithm proposed in [13, 14] could be adapted to learn stack operators without any restriction. Also, a deep neural network might also be considered to represent the characteristic function of the set W-operator and be trained by minimising the error associated with the respective stack operator. We will investigate these learning methods in the future.
6 Future research
@@
As discussed above, the next step in this research is to develop and implement learning methods for stack operators, but there are also interesting theoretical lines of research. Given the simplicity of the representation of stack W-operators, that are equivalent to Boolean functions in $\mathcal{P}(W)$, they could be the preferred choice when solving a practical image transformation problem. However, it is necessary to characterise what problems they can solve, that is, what image operators they can approximate. Therefore, the main theoretical topic for future research is to characterise the stack operators within the family of 1-Lipschitz extensions of set operators by investigating if they have universal approximating properties, i.e., by characterising spaces in which the stack operators are dense. Stack operators will be suitable to solve a problem when the optimal operator is in such a space.
Also, it would be interesting to study the representation of stack operators with other kinds of invariance, such as group equivariant operators [15] and scale invariant operators [16]. As a generalisation, other functions that map $M$ into $\{0,1\}$ besides the threshold $T_{t}$ could be considered in representation (1) yielding more general classes of grey-scale image operators. In general cases, we could obtain an operator that is an extension of a set operator, but with other Lipschitz constant and maybe with better universal approximation properties.
@@ Acknowledgements
This work was partially funded by grant #22/06211-2 São Paulo Research Foundation (FAPESP). I thank the anonymous reviewers for their comments that greatly improved the manuscript, in particular the proof of Corollary 1.
@@ Appendix 0.A Proof of results
Fix $f,g\in M^{E}$ and $\psi\in\Sigma_{S}$, and denote $\tilde{\psi}\coloneqq\mathcal{F}^{-1}(\psi)$. Define $D(f,g)=\{t\in\{1,\dots,m\}:T_{t}[f]\neq T_{t}[g]\}$ and observe that $\lVert\psi(f)-\psi(g)\rVert_{\infty}\leq\sum_{t\in D(f,g)}\lVert(\tilde{\psi}(T_{t}[f])-\tilde{\psi}(T_{t}[g]))\rVert_{\infty}\leq|D(f,g)|$. We will show that $|D(f,g)|\leq\lVert f-g\rVert_{1}$. If $x\in E$ is such that |f(x)g(x)|=s|f(x)-g(x)|=s, then $\{t_{x}+1,\dots,t_{x}+s\}\subset D(f,g)$ in which $t_{x}=\min(f(x),g(x))$ so, denoting $\Delta_{s}=\{x\in E:|f(x)-g(x)|=s\}$, we conclude that $|D(f,g)|\leq\sum_{s=1}^{m}s|\Delta_{s}|$. The result follows since $\lVert f-g\rVert_{1}=\sum_{s=1}^{m}s|\Delta_{s}|$. $\blacksquare$
@@
$=\left\{\sum_{s=1}^{m}X_{s}:X_{s}\in\mathcal{P}(W),X_{m}\leq\cdots\leq X_{1},\sum_{s=1}^{m}\mathds{1}\{X_{s}\in\mathscr{X}\}\geq t\right\}.$
@@
[…]
(1) Increasing: Direct from Proposition 3. (2) Decreasing: Analogous to increasing.
(3) Extensive: Fix $\tilde{\psi}\in\Psi^{W}$ extensive and $f\in\mathcal{P}_{M}(W)$. Then, $\psi(f)=\sum_{t=1}^{m}\tilde{\psi}(T_{t}[f])\geq\sum_{t=1}^{m}T_{t}[f]=f$ in which the inequality follows since $\tilde{\psi}$ is extensive. Fix $\psi\in\Sigma_{S}^{W}$ extensive and $X\in\mathcal{P}(W)$. Then, $\tilde{\psi}(X)=\frac{1}{m}\psi(mX)\geq\frac{1}{m}mX=X$. (4) Anti-extensive: Analogous to extensive.
(5) Erosion: Due to the main result in [2], set and image operators are erosions if, and only if, their kernel is formed by only one interval of form $[X,W]$ and $f_{t},mW]$ for all $t\geq 1$, respectively. On the one hand, it follows from Corollary 2 that if $\tilde{\psi}$ is an erosion with kernel $[X,W]$ then the kernel of $\psi$ equals $[tX,W]$ for all $t\geq 1$ so it is an erosion. On the other hand, if $\psi$ is an erosion then $\mathcal{K}_{S}(\psi)(m)=[f_{m},mW]$ and $\mathcal{K}(\tilde{\psi})=\{X\in\mathcal{P}(W):f_{m}\leq mX\}$. It is clear that $f_{m}\leq mX$ if, and only if, $T_{1}[f_{m}]\leq X$ and therefore $\mathcal{K}(\tilde{\psi})=[T_{1}[f_{m}],W]$, and $\tilde{\psi}$ is an erosion.
(6) Dilation: Analogous to erosion by considering the dual partial orders, since dilations according to a partial order are erosions according to the dual partial order.
(7) Anti-dilation: Anti-dilations are operators such that $\psi(f\vee g)=\psi(f)\wedge\psi(g)$ and $\tilde{\psi}(X\vee Y)=\tilde{\psi}(X)\wedge\tilde{\psi}(Y)$. Due to the main result in [2], the kernel of an anti-dilation is formed by a unique interval of form $\emptyset,Y]$, for set W-operators, and $\emptyset,g_{t}]$ for all $t\geq 1$, for image operators. A deduction analogous to that of erosions follows by Corollary 2.
@@
(10) Inf-generating: An inf-generating operator is the join between a dilation and an anti-erosion (ref. [2]). Analogous to sup-generating by considering the dual partial orders, since sup-generating operators according to a partial order are inf-generating according to the dual partial order.